can someone help me out with this and show me how to put a line containing the points (1,4),(2,7, and (3,10)in slope-intercept form an equation and how i would graph this i am confused
Answer by checkley71(8403)

TO CALCULATE THE SLOPE BETWEEN TWO POINTS (1,4)&(2,7) YOU MAKE A FRACTION THUS
(Y2-Y1)(X2-X1) OR
(7-4)/(2-1)=3/1=3 FOR THE SLOPE
TO VERIFY THAT THE THESE THREE POINTS ARE ON THE SAME LINE WE CALCULATE THE SLOPE OF TWO OTHER POINTS THUS
(10-4)/(3-1)=6/2=3 SAME SLOPE THUS THEY ARE ON THE SAME LINE
NOW PLOT THE THREE POINTS AND DRAW A LINE THROUGHT THEM.
THE EQUATION FOR THE LINE IS
Y=mX+b WHERE m=THE SLOPE OF 3 SO
Y=3X+b NOW SUBSTITUTE ANY OF THE THREE POINTS FOR THE Y & X VALUES IN THIS EQUATION AND SOLVE FOR THE Y INTERCEPT(b). THUS
7=3*2+b
7=6+b
b=7-6
b=1 FOR THE Y INTERCEPT (0,1)